OpenStudy (anonymous):

simplify this radical expression the cube root of 27y^6

OpenStudy (jdoe0001):

\(\Large \bf { \sqrt[3]{ 27y^6}\qquad \begin{cases} 27\to 3\cdot 3\cdot 3\to 3^3\\ y^6\to y^2\cdot y^2\cdot y^2\to (y^2)^3 \end{cases} \\ \quad \\ thus\quad \sqrt[{\color{blue}{ 3}}]{ 3^{\color{blue}{ 3}}(y^2)^{\color{blue}{ 3}}}\to ? }\)